Covert the following linear equations from y-intercept form to standard form
1.a) y=2x-3
b) y= 1/3x - 2/5
c) y= -3x+ 1/2
d) y=2x+7

Answer by Boreal(15235)

y=2x-3 is 2x-y=3
y=(1/3)x-2/5 is multiplied by 15 to make everything integers. 5x-15y=6
y=-3x+(1/2) is 6x+2y=1
y=2x+7 is 2x-y=-7